# Revisiting Data Challenges of Computational Pathology: A Pack-based Multiple Instance Learning Framework

### v1

@@ A.1 Datasets
We validate our method on various computational pathology tasks and challenging benchmarks in the ear of foundation models, including cancer grading (PANDA Bulten et al. (2022)), subtyping (TCGA-BRCA), survival analysis (TCGA-LUAD, TCGA-BRCA).
PANDA Bulten et al. (2022) is a large-scale, multi-center dataset dedicated to prostate cancer detection and grading. It comprises 10,202 digitized H&E-stained whole-slide images, making it one of the most extensive public resources for prostate cancer histopathology. Each slide is annotated according to the Gleason grading system and subsequently assigned an International Society of Urological Pathology (ISUP) grade, enabling both cancer detection and severity assessment. Specifically, ISUP Grade 1 corresponds to Gleason 3+3, Grade 2 to 3+4, Grade 3 to 4+3, Grade 4 to Gleason score 8, and Grade 5 to Gleason score 9 or 10, while Grade 0 represents benign samples. The dataset includes a diverse distribution of ISUP grades, with 2,724 slides classified as grade 0 (benign), 2,602 as grade 1, 1,321 as grade 2, 1,205 as grade 3, 1,187 as grade 4, and 1,163 as grade 5. Spanning multiple clinical centers, PANDA ensures a broad range of samples, mitigating center-specific biases.
@@
We randomly split the PANDA dataset into training, validation, and testing sets with a ratio of 7:1:2. Due to the limited data size, the remaining datasets are divided into training and testing sets with a ratio of 7:3. The grading and subtyping tasks use 5 different random seeds to ensure the stability of the results. And because the survival analysis task is more affected by data partitioning, we use 3-fold cross-validation with 3 different random seeds to conduct the experiments.
@@ A.3 Implementation Details
For our experiments conducted with a batchsize of 1, which is a conventional approach for methods handling variable sequence lengths like the two-stage methods investigated Lu et al. (2021), Shao et al. (2021), Tang et al. (2024), we consistently employed the Adam optimizer Kingma & Ba (2014). An initial learning rate of $1\times 10^{-4}$ was used, and this rate was dynamically adjusted during training using the Cosine annealing strategy. To mitigate potential overfitting and ensure robust optimization, we incorporated an early stopping mechanism across all experiments, selecting the model checkpoint that achieved the best performance on the validation metric for final evaluation. For grading tasks, training ran for a maximum of 100 epochs with a patience of 20 (starting from epoch 75). For subtyping, the maximum was 75 epochs with a patience of 20 (starting from epoch 30). For survival analysis, the maximum was 100 epochs with a patience of 10 (starting from epoch 30). To ensure a fair comparison, all baseline methods were tuned following their official guidelines and recommended hyperparameter search spaces. For experiments involving a batchsize greater than 1, the learning rate was dynamically adjusted to an empirically determined value to achieve optimal performance. Notably, specific models encountered memory limitations on the 3090 GPU when applied to certain large datasets. For instance, training the WiKG aggregator Li et al. (2024c) on the BRCA(Subtyping) dataset required sampling the number of patches down to 1024 per instance to fit into memory. Except for the aforementioned cases, all experiments were performed on NVIDIA 3090 GPUs using unified hyperparameters where applicable.
@@ Appendix B Additional Quantitative Results
B.1 More discussion about Hyperslide
Table 4: Impact of HyperSlide supervision.
Method / Task Grading Subtyping Survival
Random Sampling (RS) 77.91 93.89 65.71
RS + HyperSlide 79.93 94.37 67.15
Pack (Ours) 79.02 94.06 66.15
Pack + HyperSlide (Ours) 80.19 94.21 67.50
Effectiveness of the HyperSlide Supervision Signal. To isolate the contribution of our multi-WSI supervision mechanism, we conducted an ablation study, the results of which are presented in Tab.4. The study demonstrates that the introduction of HyperSlide labels provides a consistent and significant performance uplift across all three downstream tasks. This improvement holds true for both the baseline random sampling (RS) packing strategy and our proposed adaptive packing method. For instance, adding HyperSlide to our adaptive packing improved Grading AUC by 1.17, Subtyping AUC by 0.15, and Survival C-Index by 1.35. This underscores the efficacy of using a higher-level, clinically-informed supervision signal to guide the model’s learning process on aggregated WSI data.
Labeling Strategy / Task Grading Subtyping Survival
Mixed Soft Label 75.50 91.31 N/A
Task-specific Label (Ours) 80.19 94.21 67.50
B.2 More discussion about batched training
Strategy Grad. Sub. Surv. TTime
— Without Batched Training —
Baseline 73.21 93.58 65.87 12h
Accumulation 72.12 93.02 64.28 12h
Accumulation + patch drop 77.58 93.61 64.53 12h
Accumulation + ADS - 94.25 66.98 -
— With Batched Training —
Batched (RS) + patch drop 77.91 93.89 65.71 2h
Batched (RS) + ADS - 94.27 67.04 -
Batched (RS) + HyperSlide 79.93 94.37 67.15 2.5h
OURS 80.19 94.86 68.14 4h
Table 7: Comparison of adaptive and fixed-size packing strategies.
Strategy Grad. Sub. Surv. Pad. Ratio
Fixed-size (n=2) 79.48 94.54 67.78 54.9%
Ours (Adaptive) 80.19 94.86 68.14 18.5%
Strategy Grad. Sub. Surv.
Slide-level Mixup 76.25 93.23 N/A
Ours (HyperSlide) 80.19 94.86 67.50
B.3 More discussion about ADS
Table 9: Performance of ADS on PANDA.
Strategy CHIEF UNI GIGAP
w/ ADS 76.79 78.84 77.92
w/o ADS 76.46 80.19 80.41
Table 10: Performance of ADS on BRCA.
Strategy CHIEF UNI GIGAP
before 90.60 94.54 94.18
after 92.38 94.86 94.86
Position of the ADS Module. The position of the ADS module is a key impact of its efficacy. As shown in Tab. 10, applying the ADS module to the branched feature sets $\mathcal{R}_{b}$ and $\mathcal{D}_{b}$ independently (after) yields superior performance over applying it to the combined feature set prior to branching (before). We hypothesize that this is because operating on distinct branches enables the ADS module to learn more specialized attention mechanisms. Such specialization allows the module to better capture the unique characteristics and relevant information within each feature set ($\mathcal{R}_{b}$ and $\mathcal{D}_{b}$), which might otherwise be obscured or averaged out in a combined representation.
Table 11: Computational Cost of ADS on BRCA.
Method TTime Memory FLOPs AUC
w/o ADS 1h 7G 49.4G 94.21
w/ ADS 1.5h 13G 142.1G 94.86
Attention Mechanism. The attention mechanism is a critical component of the ADS module. In its absence, the downsampling operation would treat all instances uniformly, assigning equal importance to each. The core contribution of the attention is to compute instance-specific weights,
Method Subtyping (AUC) Survival (C-Index)
w/o Attention 94.56 67.71
w/ Attention 94.86 68.14
making the feature aggregation process both attention-driven and instance-dependent. This enables the model to selectively focus on and amplify features from the most clinically salient regions. To empirically validate its contribution, we conducted an ablation study by removing the attention component. As shown in Tab. 12, this resulted in a distinct performance degradation across both subtyping and survival prediction tasks, underscoring the mechanism’s importance in learning a meaningful data-driven downsampling policy.
@@ B.4 Ablation of Hyperparameters
B.5 Additional Benchmarking Experiments
Additional Dataset. To further substantiate the robustness and generalizability of our proposed PackMIL, we extended its evaluation to two additional, widely recognized benchmarks: a computational pathology task for cancer metastasis detection and a medical imaging task outside of pathology for diabetic retinopathy grading. The cancer metastasis detection benchmark integrates the CAMELYON-16 and CAMELYON-17 datasets Litjens et al. (2018). These datasets consist of whole-slide images (WSIs) of hematoxylin and eosin (H&E) stained lymph node sections from breast cancer patients. The primary task is to identify the presence of metastatic cancerous tissue within these lymph nodes, a critical step in cancer staging. For the non-pathology benchmark, we utilized a standard Diabetic Retinopathy (DR) Grading dataset Li et al. (2019). This dataset contains retinal fundus images and the objective is to classify them into different severity levels of diabetic retinopathy. This task serves to evaluate the model’s applicability to broader medical image classification challenges beyond histopathology. We compared PackMIL with several state-of-the-art Multiple Instance Learning (MIL) methods. As shown in Tab. 14, PackMIL demonstrates superior performance on both benchmarks. Specifically, on the CAMELYON cancer metastasis detection task, PackMIL (AB.) achieved an accuracy of 98.42%, outperforming all other methods. Similarly, in the Diabetic Retinopathy Grading task, PackMIL (AB.) obtained the highest score of 61.34%. These results underscore the effectiveness and versatility of our approach across different medical imaging domains.
Table 14: Performance comparison on additional benchmark datasets.
Method CAMELYON (UNI) Diabetic Retinopathy Grading (R50)
ABMIL 96.58 58.26
DSMIL 96.44 58.03
TransMIL 96.63 60.28
WiKG 97.08 55.84
PackMIL (AB.) 98.42 61.34
PackMIL (DS.) 97.81 60.27
Additional Encoders. To further assess the versatility and effectiveness of PackMIL, we conducted additional experiments by integrating it with more advanced feature extractors and comparing its performance against , powerful slide-level encoders. We utilized CONCHv1.5 Lu et al. (2024a) as a feature extractor and compared our model with state-of-the-art slide encoders including HIPT Chen et al. (2022b), GigaPath Xu et al. (2024), CHIEF Wang et al. (2024b), and TITAN Ding et al. (2024) on downstream tasks of tumor grading, subtyping, and survival prediction. We reported the training time (TTime) of all methods on PANDA. The results, detailed in Tab. 15, demonstrate that PackMIL consistently enhances performance while maintaining remarkable efficiency. Notably, PackMIL significantly reduces the training time, requiring only 2-3 hours, which is a substantial improvement over the 15 to 75 hours required by other encoders. This highlights PackMIL’s ability to effectively aggregate features from various powerful backbones, achieving superior predictive performance with significantly lower computational overhead.
Feature Extractor Slide Encoder Grading Subtyping Survival TTime
HIPT Chen et al. (2022b) HIPT 62.10 $\pm$ 1.06 86.93 $\pm$ 4.86 67.82 $\pm$ 9.61 32h
GigaPath Xu et al. (2024) GigaPath 65.86 $\pm$ 0.77 93.72 $\pm$ 3.42 62.64 $\pm$ 9.33 50h
GigaPath Xu et al. (2024) PackMIL 80.41 $\pm$ 0.53 94.86 $\pm$ 3.68 68.03 $\pm$ 9.10 3h
CHIEF Wang et al. (2024b) CHIEF 67.67 $\pm$ 0.84 91.43 $\pm$ 4.51 67.95 $\pm$ 8.46 15h
CONCHv1.5 Lu et al. (2024b) ABMIL 66.90 $\pm$ 0.74 95.14 $\pm$ 2.92 68.65 $\pm$ 9.17 14h
CONCHv1.5 Lu et al. (2024b) TITAN 63.72 $\pm$ 0.76 95.20 $\pm$ 2.92 - 75h
CONCHv1.5 Lu et al. (2024b) PackMIL 71.72 $\pm$ 0.65 95.43 $\pm$ 2.63 70.74 $\pm$ 7.70 2h
B.6 Inference Time Comparison
Table 16: Performance comparison of various Multiple Instance Learning (MIL) methods, restructured for improved readability on narrower page widths.
Method Inference Time (per slide) FPS
ABMIL 0.48625 ms 2056 fps
DSMIL 1.04138 ms 960 fps
TransMIL 7.05402 ms 142 fps
DTFD 3.97953 ms 251 fps
RRT-MIL 2.43291 ms 411 fps
WIKG 1.84916 ms 541 fps
PackMIL(ABMIL) 0.50414 ms 1984 fps (-4%)
PackMIL(DSMIL) 1.07524 ms 930 fps (-3%)
PackMIL(VITMIL) 1.36654 ms 731 fps
PackMIL(RRT) 2.52792 ms 396 fps (-4%)

### v2

@@ A.1 Datasets
We validate our method on various computational pathology tasks and challenging benchmarks in the ear of foundation models, including cancer grading (PANDA [panda]), subtyping (TCGA-BRCA), survival analysis (TCGA-LUAD, TCGA-BRCA).
PANDA [panda] is a large-scale, multi-center dataset dedicated to prostate cancer detection and grading. It comprises 10,202 digitized H&E-stained whole-slide images, making it one of the most extensive public resources for prostate cancer histopathology. Each slide is annotated according to the Gleason grading system and subsequently assigned an International Society of Urological Pathology (ISUP) grade, enabling both cancer detection and severity assessment. Specifically, ISUP Grade 1 corresponds to Gleason 3+3, Grade 2 to 3+4, Grade 3 to 4+3, Grade 4 to Gleason score 8, and Grade 5 to Gleason score 9 or 10, while Grade 0 represents benign samples. The dataset includes a diverse distribution of ISUP grades, with 2,724 slides classified as grade 0 (benign), 2,602 as grade 1, 1,321 as grade 2, 1,205 as grade 3, 1,187 as grade 4, and 1,163 as grade 5. Spanning multiple clinical centers, PANDA ensures a broad range of samples, mitigating center-specific biases.
@@
Table 4: Quantitative comparison between existing packing methods, Neural Image Compression with Attention (NIC-A) [aswolinskiy2025attention], and our proposed PackMIL framework across 12 benchmarks. Metrics reported are Accuracy for Grading, AUC for Sub-typing, and C-Index for Survival tasks.
Method PANDA (Grading) BRCA (Sub-typing) BRCA (Survival) LUAD (Survival)
CHIEF UNI GIGAP CHIEF UNI GIGAP CHIEF UNI GIGAP CHIEF UNI GIGAP
NIC-A [aswolinskiy2025attention] 65.48 73.21 72.91 90.22 93.65 94.13 65.36 65.87 65.72 61.99 60.54 59.88
PackMIL 76.46 80.19 80.41 93.01 94.86 94.86 69.76 70.00 68.03 64.37 63.61 63.04
Experiment Setting. We randomly split the PANDA dataset into training, validation, and testing sets with a ratio of 7:1:2. Due to the limited data size, the remaining datasets are divided into training and testing sets with a ratio of 7:3. The grading and subtyping tasks use 5 different random seeds to ensure the stability of the results. And because the survival analysis task is more affected by data partitioning, we use 3-fold cross-validation with 3 different random seeds to conduct the experiments.
@@ A.3 Implementation Details
Training Details. For our experiments conducted with a batchsize of 1, which is a conventional approach for methods handling variable sequence lengths like the two-stage methods investigated [clam, shao2021transmil, tang2024feature], we consistently employed the Adam optimizer [kingma2014adam]. An initial learning rate of $1\times 10^{-4}$ was used, and this rate was dynamically adjusted during training using the Cosine annealing strategy. To mitigate potential overfitting and ensure robust optimization, we incorporated an early stopping mechanism across all experiments, selecting the model checkpoint that achieved the best performance on the validation metric for final evaluation. For grading tasks, training ran for a maximum of 100 epochs with a patience of 20 (starting from epoch 75). For subtyping, the maximum was 75 epochs with a patience of 20 (starting from epoch 30). For survival analysis, the maximum was 100 epochs with a patience of 10 (starting from epoch 30). To ensure a fair comparison, all baseline methods were tuned following their official guidelines and recommended hyperparameter search spaces. For experiments involving a batchsize greater than 1, the learning rate was dynamically adjusted to an empirically determined value to achieve optimal performance. Notably, specific models encountered memory limitations on the 3090 GPU when applied to certain large datasets. For instance, training the WiKG aggregator [li2024dynamic] on the BRCA (Subtyping) dataset required sampling the number of patches down to 1024 per instance to fit into memory.
Training Resources. Except for the aforementioned cases, all experiments were performed on NVIDIA 3090 GPUs using unified hyperparameters where applicable.
Table 5: Details of different hyperparameters.
Hyperparameter Grad. Sub. Surv.
— PackMIL —
Downsample Ratio $k$ N/A 4 3
Hyperslide-loss Weight $\lambda$ 0.2 0.2 0.5
Branch split ratio 40% 40% 50%
ADS-pooling N/A Random Max
— Random Sampling (RS) —
Sampling Number 256 1300 1600
Hyperparameters. Tab. 5 gives the details of some important hyperparameters of PackMIL and Random Sampling (RS). Sec. B.3 gives more detailed discussion.
@@ Appendix B Additional Quantitative Results
B.1 More Discussion about Packing Method
While there have been preliminary explorations of packing strategies within computational pathology, these approaches remain nascent in scope and technical depth. Notably, methods such as Aswolinskiy et al. [aswolinskiy2025attention] rely on a coarse spatial packing technique, which merely stitches pixel-level tissue sections from the same case into a single macro-image to align with block-level labels. In stark contrast, our PackMIL framework introduces a sophisticated inter-slide-level packing mechanism that aggregates variable-length sequences from independent slides to resolve fundamental optimization bottlenecks. As evidenced by the quantitative comparisons in Table 4, our approach significantly outperforms simple concatenation strategies, demonstrating the necessity of addressing data heterogeneity through robust feature-space optimization rather than naive spatial stitching.
Feature Extractor Slide Encoder Grading Subtyping Survival TTime
HIPT [chen2022hipt] HIPT 62.10 $\pm$ 1.06 86.93 $\pm$ 4.86 67.82 $\pm$ 9.61 32h
GigaPath [gigap] GigaPath 65.86 $\pm$ 0.77 93.72 $\pm$ 3.42 62.64 $\pm$ 9.33 50h
GigaPath [gigap] PackMIL 80.41 $\pm$ 0.53 94.86 $\pm$ 3.68 68.03 $\pm$ 9.10 3h
CHIEF [chief] CHIEF 67.67 $\pm$ 0.84 91.43 $\pm$ 4.51 67.95 $\pm$ 8.46 15h
CONCHv1.5 [lu2024visual] ABMIL 66.90 $\pm$ 0.74 95.14 $\pm$ 2.92 68.65 $\pm$ 9.17 14h
CONCHv1.5 [lu2024visual] TITAN 63.72 $\pm$ 0.76 95.20 $\pm$ 2.92 - 75h
CONCHv1.5 [lu2024visual] PackMIL 71.72 $\pm$ 0.65 95.43 $\pm$ 2.63 70.74 $\pm$ 7.70 2h
B.2 Additional Benchmarking Experiments
Additional Dataset. To further substantiate the robustness and generalizability of our proposed PackMIL, we extended its evaluation to two additional, widely recognized benchmarks: a computational pathology task for cancer metastasis detection and a medical imaging task outside of pathology for diabetic retinopathy grading. The cancer metastasis detection benchmark integrates the CAMELYON-16 and CAMELYON-17 datasets [litjens20181399]. These datasets consist of whole-slide images (WSIs) of hematoxylin and eosin (H&E) stained lymph node sections from breast cancer patients. The primary task is to identify the presence of metastatic cancerous tissue within these lymph nodes, a critical step in cancer staging. For the non-pathology benchmark, we utilized a standard Diabetic Retinopathy (DR) Grading dataset [li2019diagnostic]. This dataset contains retinal fundus images and the objective is to classify them into different severity levels of diabetic retinopathy. This task serves to evaluate the model’s applicability to broader medical image classification challenges beyond histopathology. We compared PackMIL with several state-of-the-art Multiple Instance Learning (MIL) methods. As shown in Tab. 7, PackMIL demonstrates superior performance on both benchmarks. Specifically, on the CAMELYON cancer metastasis detection task, PackMIL (AB.) achieved an accuracy of 98.42%, outperforming all other methods. Similarly, in the Diabetic Retinopathy Grading task, PackMIL (AB.) obtained the highest score of 61.34%. These results underscore the effectiveness and versatility of our approach across different medical imaging domains.
Table 7: Performance comparison on additional benchmark datasets.
Method CAMELYON (UNI) DR (R50)
ABMIL 96.58 58.26
DSMIL 96.44 58.03
TransMIL 96.63 60.28
WiKG 97.08 55.84
PackMIL (AB.) 98.42 61.34
PackMIL (DS.) 97.81 60.27
Additional Encoders. To further assess the versatility and effectiveness of PackMIL, we conducted additional experiments by integrating it with more advanced feature extractors and comparing its performance against , powerful slide-level encoders. We utilized CONCHv1.5 [lu2024avisionlanguage] as a feature extractor and compared our model with state-of-the-art slide encoders including HIPT [chen2022hipt], GigaPath [gigap], CHIEF [chief], and TITAN [ding2024titan] on downstream tasks of tumor grading, subtyping, and survival prediction. We reported the training time (TTime) of all methods on PANDA. The results, detailed in Tab. 6, demonstrate that PackMIL consistently enhances performance while maintaining remarkable efficiency. Notably, PackMIL significantly reduces the training time, requiring only 2-3 hours, which is a substantial improvement over the 15 to 75 hours required by other encoders. This highlights PackMIL’s ability to effectively aggregate features from various powerful backbones, achieving superior predictive performance with significantly lower computational overhead.
@@ B.3 Ablation of Hyperparameters
B.4 More Discussion about Hyperslide
Effectiveness of the HyperSlide Supervision. To isolate the contribution of our multi-WSI supervision mechanism, we conducted an ablation study, the results of which are presented in Tab. 9. The study demonstrates that the introduction of HyperSlide labels provides a consistent and significant performance uplift across all three downstream tasks. This improvement holds true for both the baseline random sampling (RS) packing strategy and our proposed adaptive packing method. For instance, adding HyperSlide to our adaptive packing improved Grading AUC by 1.17, Subtyping AUC by 0.15, and Survival C-Index by 1.35. This underscores the efficacy of using a higher-level, clinically-informed supervision signal to guide the model’s learning process on aggregated WSI data.
Table 9: Impact of HyperSlide supervision.
Method / Task Grading Subtyping Survival
Random Sampling (RS) 77.91 93.89 65.71
RS + HyperSlide 79.93 94.02 67.15
Pack (Ours) 79.02 94.06 66.15
Pack + HyperSlide (Ours) 80.19 94.21 67.50
Labeling Strategy / Task Grading Subtyping Survival
Mixed Soft Label 75.50 91.31 N/A
Task-specific Label (Ours) 80.19 94.21 67.50
Strategy Grad. Sub. Surv. TTime
— Without Batched Training —
Baseline 73.21 93.58 65.87 12h
Accumulation 72.12 93.02 64.28 12h
Accumulation + patch drop 77.58 93.61 64.53 12h
Accumulation + ADS - 94.25 66.98 -
— With Batched Training —
Batched (RS) + patch drop 77.91 93.89 65.71 2h
Batched (RS) + ADS - 94.27 67.04 -
Batched (RS) + HyperSlide 79.93 94.02 67.15 2.5h
Ours 80.19 94.86 68.14 4h
B.5 More Discussion about Batched Training
Table 12: Comparison of adaptive and fixed-size packing strategies.
Strategy Grad. Sub. Surv. Pad. Ratio
Fixed-size (n=2) 79.48 94.54 67.78 54.9%
Ours (Adaptive) 80.19 94.86 68.14 18.5%
Strategy Grad. Sub. Surv.
Slide-level Mixup 76.25 93.23 N/A
Ours (HyperSlide) 80.19 94.86 67.50
B.6 More Discussion about ADS
Table 14: Performance of ADS on PANDA.
Strategy CHIEF UNI GIGAP
w/ ADS 76.79 78.84 77.92
w/o ADS 76.46 80.19 80.41
Figure 6: Attention visualization on the PANDA dataset [panda]. The RS strategy, due to its sampling, exhibits limited global attention. With multi-slide supervision via hyperslides and the supplementation of key features, PackMIL demonstrates a more accuracy and comprehensive focus on tumor areas.
Position of the ADS Module. The position of the ADS module is a key impact of its efficacy. As shown in Tab. 15, applying the ADS module to the branched feature sets $\mathcal{K}_{b}$ and $\mathcal{D}_{b}$ independently (after) yields superior performance over applying it to the combined feature set prior to branching (before). We hypothesize that this is because operating on distinct branches enables the ADS module to learn more specialized attention mechanisms. Such specialization allows the module to better capture the unique characteristics and relevant information within each feature set ($\mathcal{K}_{b}$ and $\mathcal{D}_{b}$), which might otherwise be obscured or averaged out in a combined representation.
Table 15: Performance of ADS on BRCA.
Strategy CHIEF UNI GIGAP
before 90.60 94.54 94.18
after 92.38 94.86 94.86
Table 16: Computational Cost of ADS on BRCA.
Method TTime Memory FLOPs AUC
w/o ADS 1h 7G 49.4G 94.21
w/ ADS 1.5h 13G 142.1G 94.86
Attention Mechanism. The attention mechanism is a critical component of the ADS module. In its absence, the downsampling operation would treat all instances uniformly, assigning equal importance to each. The core contribution of the attention is to compute instance-specific weights, making the feature aggregation process both attention-driven and instance-dependent. This enables the model to selectively focus on and amplify features from the most clinically salient regions. To empirically validate its contribution, we conducted an ablation study by removing the attention component. As shown in Tab. 17, this resulted in a distinct performance degradation across both subtyping and survival prediction tasks, underscoring the mechanism’s importance in learning a meaningful data-driven downsampling policy.
Method Subtyping (AUC) Survival (C-Index)
w/o Attention 94.56 67.71
w/ Attention 94.86 68.14
B.7 Inference Time Comparison
Table 18: Inference Efficiency comparison of various MIL methods. Since PackMIL only adds the ADS module during inference and retains the simplest MIL inference pipeline, its inference time is nearly identical to the baseline.
Method Inference Time FPS
ABMIL 0.48 ms 2056 fps
DSMIL 1.04 ms 960 fps
TransMIL 7.05 ms 142 fps
DTFD 3.97 ms 251 fps
RRT-MIL 2.43 ms 411 fps
WIKG 1.84 ms 541 fps
PackMIL(ABMIL) 0.50 ms 1984 fps (-4%)
PackMIL(DSMIL) 1.07 ms 930 fps (-3%)
PackMIL(VITMIL) 1.36 ms 731 fps
PackMIL(RRT) 2.52 ms 396 fps (-4%)